| <p>This morning at 8:35am the pressures read (in mBar):</p>
<ul>
<li>TPU1 = 2.0e-6</li>
<li>TPU2 = 4.5e-7</li>
<li>TPD1 = 8.1e-7</li>
<li>TPD2 = 7.9e-7</li>
<li>CHUP = 4.7e-7</li>
<li>CHDW = 1.1e-5</li>
</ul>
<p>The detectors were biased. Far S2 leak I = 1.54uA, near S2 leak I = 12.53uA.</p>
<p>It was discovered the ECL-NIM-ECL converter located immediately to the left of the pulser has broken channels - the last 8 channels. Instead of connecting the 32 pin ribbon cable to the input and taking lemo outputs 5 and 13, we are now connecting two single channel pin cables to inputs 1 and 2, and taking the corresponding lemo outputs.</p>
<p>We found RAL X sometimes doesn't produce a trigger for the pulser.</p>
<p>We replaced RAL I with RAL XIV, RAL XIV is bad.</p>
<p>We replaced RAL XIV with RAL XV, RAL XV is good.</p>
<p>We finished setting the RAL thresholds.</p> |